OpenStudy (anonymous):

According to the rational root theorem, what are possible roots of the polynomial function below? F(x) = 8x3 - 3x2 + 5x + 15

OpenStudy (anonymous):

there are no rational roots to that function

OpenStudy (anonymous):

really? because:

OpenStudy (anonymous):

ask the teacher?

OpenStudy (anonymous):

can't

OpenStudy (anonymous):

i have to take these quizzes at home

OpenStudy (anonymous):

@superhelp101

OpenStudy (anonymous):

oh. u dont need to select any

OpenStudy (anonymous):

it says "select all correct answers" and none are correcet

OpenStudy (anonymous):

yeah i do..it won't let me not select any

OpenStudy (anonymous):

that's why i'm confused

OpenStudy (anonymous):

ihni then. sorry :/

OpenStudy (anonymous):

it's ok

OpenStudy (anonymous):

@Melodysim

OpenStudy (anonymous):

@Here_to_Help15 do you know this?

OpenStudy (aum):

F(x) = 8x3 - 3x2 + 5x + 15 Possible rational roots are: \(\Large \pm \frac{\text{factors of 15}}{\text{factors of 8}} \)

OpenStudy (anonymous):

soooo..like 5/2?

OpenStudy (anonymous):

-1/4?

OpenStudy (aum):

factors of 15: 1, 3, 5, 15 factors of 8: 1, 2, 4, 8 Possible rational roots are: \(\Large \pm \frac{1,3,5,15}{1,2,4,8} \)

OpenStudy (anonymous):

so A and E (is that all?)

OpenStudy (aum):

\(\Large \pm \frac{1,3,5,15}{1,2,4,8}\) Go through each answer choice and see if it is one of the choices from above. 5/2 - yes 4 - no 5 - yes (5/1) -3 - yes (-3/1) -1/4 - yes 2/3 - no

OpenStudy (anonymous):

THANKS!

OpenStudy (aum):

yw.